# Evidence for a Sharp CO Snowline Transition in a Protoplanetary Disk and Implications for Millimeter-wave Observations of CO Isotopologues

### v1

@@ Abstract
Observations of CO isotopologue emission from protoplanetary disks at millimeter wavelengths are a powerful tool for probing the CO snowline, an important marker for disk chemistry, and also for estimating total disk gas mass, a key quantity for planet formation. We use simple models to demonstrate that the vertical thickness of an isothermal layer around the disk midplane has important effects on the CO column density radial profile, with a thick layer producing a sharp CO snowline transition. We simulate ngVLA and ALMA images to show that this sharp change in CO column density can be detected in the derivative of the radial profile of emission from optically thin CO isotopologue lines. We apply this method to archival ALMA observations of the disk around the Herbig Ae star HD 163296 in the $\mathrm{C^{17}O}$ and $\mathrm{C^{18}O}$ J=1-0 and J=2-1 lines to identify a sharp CO snowline transition near $\sim 80$ au ($0\farcs 8$ at 101 pc), and show the CO column density decreases by more than a factor of 20. This finding is consistent with previous inferences from the steep rise of N2H+ emission, which marks the location where CO depletes. We also demonstrate that the disk’s thermal structure introduces a significant systematic uncertainty to estimates of total disk gas mass derived from these lines. The substantial improvement in sensitivity envisioned for the ngVLA over ALMA for observations of ground state lines of CO isotopologues has the potential to extend this approach to a much larger population of disks.
@@ 1 Introduction
Planets form from dust and gas within protoplanetary disks surrounding young stars. In the cold disk midplane, temperatures are low enough that many molecules are frozen out onto grain mantles and depleted from the gas phase. The CO snowline, defined as the radius at which the midplane temperature is low enough for CO to freeze out onto dust grains, is an important feature of these disks. The CO snowline location is pivotal because of its effects on the chemical composition and physical properties of the disk. The locations of disk substructures relative to the snowlines (“condensation fronts”) of major volatile species, including H2O, CO2, CO, and N2, are important for setting the chemistry in the disks (Öberg et al., 2023). Radial variations in the C/O ratio affect the composition of material accreted by forming planets, influencing the abundances in their atmospheres (e.g., Öberg et al., 2011). In addition, CO, and its less abundant isotopologues, are also commonly used as tracers of disk gas mass, essential for elucidating processes involved in planet building, as cold H2 – the primary mass constituent – is impossible to observe directly. However, an accurate conversion of CO emission to total gas mass requires accounting for CO that has frozen out of the gas phase, as well as photodissociation and potential chemical conversion processes (Ansdell et al., 2018; Bosman et al., 2018; Krijt et al., 2018).
@@
The cold midplane and its proximity to chemically active layers in disks can be probed directly by observations of low excitation lines of trace species at millimeter wavelengths. The CO snowline is difficult to observe directly because of radial and vertical temperature gradients, leading to a warmer region above the “snow surface” where CO remains in the gaseous state throughout (Qi et al., 2011). While emission from the most abundant main isotopologue 12CO in disks is bright, it is also generally very optically thick and traces only an effective photospheric temperature high above the midplane. By contrast, emission from lower abundance CO isotopologues (13CO, C18O, C17O, 13C18O) have lower optical depths and emerge from deeper in the disk. The high sensitivity and resolution of ALMA for observing low-J lines of these rarer CO isotopologues, and other trace species of cold gas in disks, has revolutionized studies of CO snowlines (e.g., Zhang et al., 2017; Booth et al., 2023). Images of N2H+, a species whose abundance is greatly enhanced where freeze-out removes CO from the gas phase (Bergin et al., 2002) has shown promise as a chemical tracer of the CO snowline (Qi et al., 2013), although the interpretation is complicated by the three dimensional structure of the condensation front that arises from the details of the vertical temperature structure (van ’t Hoff et al., 2017; Qi et al., 2019). More directly, high resolution observations of CO isotopologues, together with disk models, have been used to define the three dimensional CO “snow surface” in disks, though this method is complicated by the fact that the most easily detected CO isotopologue lines can have substantial opacity (e.g, Pinte et al., 2018).
@@
The temperature structure of protoplanetary disks is set by a balance of heating and cooling that depends sensitively on the effects grain growth, and size-dependent radial drift and vertical settling (e.g., D’Alessio et al., 2006). The interplay of these processes can have important effects on the location of the CO freeze-out region. For example, the models of Cleeves (2016) predict that inward drift of millimeter-sized grains can result in a small radial temperature inversion that creates a second CO snowline. A more significant effect is predicted to stem from the relative scale heights of large and small grains that controls the extent of a Vertically Isothermal Region around the Midplane (VIRaM) layer, where temperatures are nearly isothermal vertically at the CO freeze-out region. Qi et al. (2019) showed that disks with a thick VIRaM layer exhibit bright, narrow N2H+ rings that peak in the slim radial zone between the CO and N2 snowlines. While this chemical signature of a thick VIRaM layer is highly suggestive, direct evidence is still lacking.
@@
In this paper, Section 2 presents a generic disk model to show that a thick VIRaM layer leads to a sharp drop in CO column density across the CO snowline, by more than an order of magnitude, and that direct detection of such a “sharp CO snowline transition” can be made using observations of low excitation rotational lines of rare CO isotopologues with high sensitivity, spatial, and spectral resolution. Section 3 applies the concepts from this model to archival ALMA observations of CO isotopologues from the large disk around the nearby Herbig Ae star HD 163296, hypothesized to possess a thick VIRaM layer from its N2H+ emission morphology, and provides strong evidence for the presence of a sharp CO snowline transition. Section 4 discusses implications of this thermal structure, in particular for disk gas mass estimates obtained from observations of CO isotopologues. In this context, the next generation Very Large Array (ngVLA, Murphy et al., 2018) has potential for substantially advancing this field as a result of its substantially higher sensitivity than ALMA for observations of rare CO isotopologue ground state lines.
@@ 2 Disk Vertical Thermal Structure Effects
To investigate the impact of the disk vertical thermal structure on the CO radial column density profile, we employ a simple parametric disk model similar to that used by Andrews et al. (2009). We then simulate observations of the J=1–0 and J=2–1 line emission of rare CO isotopologues to demonstrate the observable consequences of the VIRaM layer thickness.
@@ 2.1 Disk Model Prescriptions
Figure 1: Disk densities and temperatures for two models: (left) thick VIRaM layer with an isothermal vertical temperature structure, (center) thin VIRaM layer with a conventional vertical temperature gradient. Temperature contours (red) are at 20, 24, and 28 K, and density contours (blue) are at $10^{8}$, $10^{9}$, and $10^{10}$ cm-3. (right) The CO column density distributions for the two models. The disk with a thick VIRaM layer shows a sharp radial drop across the CO snowline.
We assume CO emission originates in a vertical layer of the disk with a constant abundance $f_{\mathrm{CO}}=1\times 10^{-4}$. The upper boundary of this layer, toward the disk surface, is constrained by CO photodissociation, which occurs above an H2 surface column density of approximately $10^{21}\,\text{cm}^{-2}$ (Qi et al., 2011). The lower boundary, closer to the midplane, is defined by the CO freeze-out temperature, assumed to be 24 K. When the temperature drops below 24 K, the depletion factor – the reduction in CO fractional abundance due to freeze-out – is set to 320. In such regions, the CO abundance decreases to $3.1\times 10^{-7}$. The right panel of Figure 1 displays the resulting CO column density profiles for both models. The model with a thick VIRaM layer exhibits a very sharp radial drop in CO column density at the snowline (80 au), whereas the model with a thin VIRaM layer shows a more gradual decrease.
These models serve to illustrate that the disk vertical thermal structure significantly impacts the CO radial column density profile. For the thick VIRaM layer model, temperature contours in the disk region below the CO freeze-out temperature are nearly perpendicular to the density contours, which produces a very sharp drop in CO column density across the snowline. In contrast, the thin VIRaM layer model shows a much smoother gradient in CO column density.
@@ 2.2 CO Isotopologue Imaging Simulations
The CO snowline can be probed best using optically thin tracers sensitive to cold gas, such as the lowest excitation rotational lines of rare CO isotopologues, together with sufficiently high spatial resolution — specifically, a beam size smaller than half the CO snowline radius — to resolve the column density change within the beam. Low optical depth provides intrinsic sensitivity to changes in CO column density with radius, while high spatial resolution prevents the shape of the radial profile from being smoothed out.
Using the CO distributions from two models with thin and thick VIRaM layers and assuming local interstellar medium ratios of CO/C17O = 2005 and CO/C18O = 557 (Wilson, 1999), we simulated $\mathrm{C^{17}O}$ and $\mathrm{C^{18}O}$ 1-0 and 2-1 emission using the non-local thermodynamic equilibrium, two-dimensional accelerated Monte Carlo radiative transfer code RATRAN (Hogerheijde & van der Tak, 2000). The outputs of these simulations were then converted into synthetic visibility datasets using the simobserve task in the Common Astronomy Software Application package (CASA Team et al., 2022).
@@
For the weaker $\mathrm{C^{17}O}$ and $\mathrm{C^{18}O}$ 1-0 lines, we used the ngVLA for the imaging simulations due to its superior sensitivity at these frequencies compared to ALMA. The ngVLA simulations used the configuration file ngvla-revD.core.cfg, which includes 114 antennas with a diameter of 18 meters within a 4 km region, providing a beam size of about $0\farcs 3$, comparable to that for the ALMA simulations. The simulated observations had an on-source integration time of 4 hours for ngVLA simulations. To add proper noise to the ngVLA simulations, we followed the recommendations from the CASA simulation guide [^1] https://casaguides.nrao.edu/index.php/Simulating_ngVLA_Data-CASA5.4.1 and estimated the scaling parameter $\sigma\_{simple}$, then corrupted the simulated data using the sm.setnoise function of the sm toolkit. Note that ALMA would require an integration time of about 30 times longer than ngVLA to reach the same noise level for these lines at this resolution, according to the available on-line telescope sensitivity estimators for ALMA [^2] https://almascience.nrao.edu/proposing/sensitivity-calculator (for typical band 3 weather) and ngVLA [^3] https://ngect.nrao.edu/.
The detailed parameters for each simulated line are summarized in Table 2. The source declination was assumed to be $-22$ degrees. The disk was assumed to be inclined by $45$ degrees oriented at a position angle $90$ degrees (east from north). The simulations were conducted with 60 velocity channels of width 0.5 km -1 for each line. Imaging was performed using the tclean task in CASA with Briggs weighting and a robust parameter of $2$, i.e. close to natural weighting for maximum sensitivity. Finally we collapsed the spectral cubes into integrated intensity maps using the Python package bettermoments (Teague & Foreman-Mackey, 2018).
@@
$\mathrm{C^{17}O}$ 1-0 112.359 0.5 14400 ngvla-revD.core.cfg $0.32\times 0.20$ 0.24
@@
$\mathrm{C^{18}O}$ 1-0 109.782 0.5 14400 ngvla-revD.core.cfg $0.32\times 0.21$ 0.18
@@
We extracted the deprojected and azimuthally averaged radial intensity profiles using the radial_profile function from the Python package GoFish (Teague, 2019). Figure 3 (top panels) shows the radial profiles of the integrated intensity of the simulated C17O and C18O 1-0 and 2-1 emission for models with thick and thin VIRaM layers. While a sharp drop in CO column density with radius is not immediately evident from the radial profiles of the integrated emission intensity, it can be identified by analyzing the derivative of these radial profile. If the disk has a thick VIRaM layer, then the resulting sharp CO snowline will present a local minimum in the derivative, provided the emission is optically thin. Figure 3 (bottom panels) show the derivatives of the radial profiles, with the location of the CO snowline indicated by a gray dashed line.
The derivatives for the thick VIRaM layer model show a clear minimum near the CO snowline for the J=1–0 lines, marking a significant change in the rate of CO column density decline with radius. In contrast, the derivatives for the thin VIRaM layer model do not show a minimum clearly associated with the CO snowline. The minimum is less prominent in the C17O 2-1 profile and is not visible at all in the C18O 2-1 profile, as the emission from this more abundance isotopologue is optically thick throughout. This highlights that the “sharpness” of the CO snowline transition is best distinguished using the most optically thin tracers, and particularly the ground state lines. This radial profile derivative analysis provides an empirical way to determine the CO snowline location.
Figure 3: Radial profiles of integrated intensity (top) and their derivatives (bottom) for simulated C17O 1-0, C18O 1-0 C17O 2-1, C18O 2-1 emission (from left to right) in models with thick (blue) and thin (orange) VIRaM layers. The grey dashed line in each panel indicates the location of the CO snowline. For the optically thin lines, the derivative profiles for the thick VIRaM model show a minimum near the CO snowline resulting from the significant change in CO column density with radius. The minimum in the derivative is less prominent in C18O 1-0 than in C17O 1-0 due to increasing optical depth, and it becomes undetectable in C18O 2-1, which is optically thick.

### v2

@@ Abstract
Observations of CO isotopologue emission from protoplanetary disks at millimeter wavelengths are a powerful tool for probing the CO snowline, an important marker for disk chemistry, and also for estimating total disk gas mass, a key quantity for planet formation. We use simple models to demonstrate that the vertical thickness of an isothermal layer around the disk midplane has important effects on the CO column density radial profile, with a thick layer producing a sharp CO snowline transition. We simulate ngVLA and ALMA images to show that this sharp change in CO column density can be detected in the derivative of the radial profile of emission from optically thin CO isotopologue lines. We apply this method to archival ALMA observations of the disk around the Herbig Ae star HD 163296 in the $\mathrm{C^{17}O}$ and $\mathrm{C^{18}O}$ J=1-0 and J=2-1 lines to identify a sharp CO snowline transition near $\sim 80$ au ($0\farcs 8$ at 101 pc), and show the CO column density decreases by more than a factor of 20. This finding is consistent with previous inferences from the steep rise of N2H+ emission, which marks the location where CO depletes. We also demonstrate that the disk’s thermal structure introduces significant systematic uncertainty to estimates of total disk gas mass derived from these lines. The substantial improvement in sensitivity envisioned for the ngVLA over ALMA for observations of ground-state lines of CO isotopologues has the potential to extend this approach to a much larger population of disks.
@@ 1 Introduction
Planets form from dust and gas within protoplanetary disks surrounding young stars. In the cold disk midplane, temperatures are low enough that many molecules are frozen out onto grain mantles and depleted from the gas phase. The CO snowline, defined as the radius at which the midplane temperature is low enough for CO to freeze out onto dust grains, is an important feature of these disks. The CO snowline location is pivotal because of its effects on the chemical composition and physical properties of the disk. The locations of disk substructures relative to the snowlines (“condensation fronts”) of major volatile species, including H2O, CO2, CO, and N2, are important for setting the chemistry in the disks (Öberg et al., 2023). Radial variations in the C/O ratio affect the composition of material accreted by forming planets, influencing the abundances in their atmospheres (e.g., Öberg et al., 2011). In addition, CO, and its less abundant isotopologues, are also commonly used as tracers of disk gas mass, essential for elucidating processes involved in planet building, as cold H2 – the primary mass constituent – is impossible to observe directly. However, accurate conversion of CO emission to total gas mass requires accounting for CO that has frozen out of the gas phase, as well as photodissociation and potential chemical conversion processes (Ansdell et al., 2018; Bosman et al., 2018; Krijt et al., 2018).
@@
The cold midplane and its proximity to chemically active layers in disks can be probed directly by observations of low excitation lines of trace species at millimeter wavelengths. The CO snowline is difficult to observe directly because of radial and vertical temperature gradients, leading to a warmer region above the “snow surface” where CO remains in the gaseous state throughout (Qi et al., 2011). While emission from the most abundant main isotopologue 12CO in disks is bright, it is also generally very optically thick and traces only an effective photospheric temperature high above the midplane. By contrast, emission from lower-abundance CO isotopologues (13CO, C18O, C17O, 13C18O) have lower optical depths and emerge from deeper in the disk. The high sensitivity and resolution of ALMA for observing low-$J$ lines of these rarer CO isotopologues, and other trace species of cold gas in disks, has revolutionized studies of CO snowlines (e.g., Zhang et al., 2017; Booth et al., 2023). N2H+, a species whose abundance is greatly enhanced where freeze-out removes CO from the gas phase (Bergin et al., 2002) has shown promise as a chemical tracer of the CO snowline (Qi et al., 2013), although the interpretation of N2H+is complicated by the three-dimensional structure of the condensation front that arises from the details of the vertical temperature structure (van ’t Hoff et al., 2017; Qi et al., 2019). More directly, high-resolution observations of CO isotopologues, together with disk models, have been used to define the three-dimensional CO “snow surface” in disks, though this method is complicated by the fact that the most easily detected CO isotopologue lines can have substantial opacity (e.g, Pinte et al., 2018).
@@
The temperature structure of protoplanetary disks is set by a balance of heating and cooling that depends sensitively on the effects of grain growth, and size-dependent radial drift and vertical settling (e.g., D’Alessio et al., 2006). The interplay of these processes can have important effects on the location of the CO freeze-out region. For example, the models of Cleeves (2016) predict that inward drift of millimeter-sized grains can result in a small radial temperature inversion that creates a second CO snowline. A more significant effect is predicted to stem from the relative scale heights of large and small grains that control the extent of a vertically isothermal region around the midplane (VIRaM) layer, where temperatures are nearly isothermal vertically at the CO freeze-out region. Qi et al. (2019) showed that disks with a thick VIRaM layer exhibit bright, narrow N2H+ rings that peak in the slim radial zone between the CO and N2 snowlines. While this chemical signature of a thick VIRaM layer is highly suggestive, direct evidence is still lacking.
@@
In this paper, Section 2 presents a generic disk model to show that a thick VIRaM layer leads to a sharp drop in the CO column density across the CO snowline, by more than an order of magnitude, and that direct detection of such a “sharp CO snowline transition” can be made using observations of low excitation rotational lines of rare CO isotopologues with high sensitivity, spatial resolution, and spectral resolution. Section 3 applies the concepts from this model to archival ALMA observations of CO isotopologues from the large disk around the nearby Herbig Ae star HD 163296, hypothesized to possess a thick VIRaM layer from its N2H+ emission morphology, and provides strong evidence for the presence of a sharp CO snowline transition. Section 4 discusses the implications of this thermal structure, in particular for disk gas mass estimates obtained from observations of CO isotopologues. In this context, the next generation Very Large Array (ngVLA, Murphy et al., 2018) has the potential for substantially advancing this field as a result of its substantially higher sensitivity than ALMA for observations of rare CO isotopologue ground-state lines.
@@ 2 Disk Vertical Thermal Structure Effects
To investigate the impact of the disk’s vertical thermal structure on the CO radial column density profile, we employ a simple parametric disk model similar to that used by Andrews et al. (2009). We then simulate observations of the J=1–0 and J=2–1 line emission of rare CO isotopologues to demonstrate the observable consequences of the VIRaM layer thickness.
@@ 2.1 Disk Model Prescriptions
Figure 1: Disk densities and temperatures for two models: (left) thick VIRaM layer with an isothermal vertical temperature structure, (center) thin VIRaM layer with a conventional vertical temperature gradient. The temperature contours (red) are at 20, 24, and 28 K, and the density contours (blue) are at $10^{8}$, $10^{9}$, and $10^{10}$ cm-3. (right) CO column density distributions for the two models. The disk with a thick VIRaM layer shows a sharp radial drop across the CO snowline.
We assume CO emission originates in a vertical layer of the disk with a constant abundance $f_{\mathrm{CO}}=1\times 10^{-4}$. The upper boundary of this layer, toward the disk surface, is constrained by CO photodissociation, which occurs above an H2 surface column density of approximately $10^{21}\,\text{cm}^{-2}$ (Qi et al., 2011). The lower boundary, closer to the midplane, is defined by the CO freeze-out temperature, assumed to be 24 K. When the temperature drops below 24 K, the depletion factor -– the reduction in CO fractional abundance due to freeze-out -– is set to 320. In such regions, the CO abundance decreases to $3.1\times 10^{-7}$. The right panel of Figure 1 displays the resulting CO column density profiles for both models. The model with a thick VIRaM layer exhibits a very sharp radial drop in CO column density at the snowline (80 au), whereas the model with a thin VIRaM layer shows a more gradual decrease.
These models serve to illustrate that the disk’s vertical thermal structure significantly impacts the CO radial column density profile. For the thick VIRaM layer model, temperature contours in the disk region below the CO freeze-out temperature are nearly perpendicular to the density contours, which produces a very sharp drop in the CO column density across the snowline. In contrast, the thin VIRaM layer model shows a much smoother gradient in CO column density.
@@ 2.2 CO Isotopologue Imaging Simulations
The CO snowline can be probed best using optically thin tracers sensitive to cold gas, such as the lowest excitation rotational lines of rare CO isotopologues, together with sufficiently high spatial resolution – in particular, a beam size smaller than half the CO snowline radius – to resolve the column density change within the beam. Low optical depth provides intrinsic sensitivity to changes in the CO column density with the radius, while high spatial resolution prevents the shape of the radial profile from being smoothed out.
Using the CO distributions from two models with thin and thick VIRaM layers and assuming local interstellar medium (ISM) ratios of CO/C17O = 2005 and CO/C18O = 557 (Wilson, 1999), we simulated $\mathrm{C^{17}O}$ and $\mathrm{C^{18}O}$ 1-0 and 2-1 emission using the nonlocal thermodynamic equilibrium, two-dimensional accelerated Monte Carlo radiative transfer code RATRAN (Hogerheijde & van der Tak, 2000). The outputs of these simulations were then converted into synthetic visibility datasets using the simobserve task in the Common Astronomy Software Application package (CASA Team et al., 2022).
@@
For the weaker $\mathrm{C^{17}O}$ and $\mathrm{C^{18}O}$ 1-0 lines, we used the ngVLA for the imaging simulations due to its superior sensitivity at these frequencies compared to ALMA. The ngVLA simulations used the configuration file ngvla-revD.core.cfg, which includes 114 antennas with a diameter of 18 meters within a 4 km region, providing a beam size of about $0\farcs 3$, comparable to that for the ALMA simulations. The simulated observations had an on-source integration time of 4 hours for ngVLA simulations. To add proper noise to the ngVLA simulations, we followed the recommendations from the CASA simulation guide [^1] https://casaguides.nrao.edu/index.php/Simulating_ngVLA_Data-CASA5.4.1 and estimated the scaling parameter $\sigma\_{simple}$, then corrupted the simulated data using the sm.setnoise function of the sm toolkit. Note that ALMA would require an integration time of about 30 times longer than ngVLA to reach the same noise level for these lines at this resolution, according to the available online telescope sensitivity estimators for ALMA [^2] https://almascience.nrao.edu/proposing/sensitivity-calculator (for typical band 3 weather) and ngVLA [^3] https://ngect.nrao.edu/.
The detailed parameters for each simulated line are summarized in Table 2. The source declination was assumed to be $-22$ degrees. The disk was assumed to be inclined by $45$ degrees oriented at a position angle of $90$ degrees (east from north). The simulations were conducted with 60 velocity channels of width 0.5 km -1 for each line. Imaging was performed using the tclean task in CASA with Briggs weighting and a robust parameter of $2$, i.e. close to natural weighting for maximum sensitivity. Finally, we collapsed the spectral cubes into integrated intensity maps using the Python package bettermoments (Teague & Foreman-Mackey, 2018).
@@
$\mathrm{C^{17}O}$ 1-0 112.359 0.5 14,400 ngvla-revD.core.cfg $0.32\times 0.20$ 0.24
@@
$\mathrm{C^{18}O}$ 1-0 109.782 0.5 14,400 ngvla-revD.core.cfg $0.32\times 0.21$ 0.18
@@
We extracted the deprojected and azimuthally averaged radial intensity profiles using the radial_profile function from the Python package GoFish (Teague, 2019). Figure 3 (top panels) shows the radial profiles of the integrated intensity of the simulated C17O and C18O 1-0 and 2-1 emission for models with thick and thin VIRaM layers. While a sharp drop in the CO column density with the radius is not immediately evident from the radial profiles of the integrated emission intensity, it can be identified by analyzing the derivative of this radial profile. If the disk has a thick VIRaM layer, then the resulting sharp CO snowline will present a local minimum in the derivative, provided the emission is optically thin. Figure 3 (bottom panels) shows the derivatives of the radial profiles, with the location of the CO snowline indicated by a gray dashed line.
The derivatives for the thick VIRaM layer model show a clear minimum near the CO snowline for the J=1–0 lines, marking a significant change in the rate of CO column density decline with the radius. In contrast, the derivatives for the thin VIRaM layer model do not show a minimum clearly associated with the CO snowline. The minimum is less prominent in the C17O 2-1 profile and is not visible at all in the C18O 2-1 profile, as the emission from this more abundant isotopologue is optically thick throughout. This highlights that the “sharpness” of the CO snowline transition is best distinguished using the most optically thin tracers, particularly the ground-state lines. This radial profile derivative analysis provides an empirical way to determine the CO snowline location.
Figure 3: Radial profiles of integrated intensity (top) and their derivatives (bottom) for simulated C17O 1-0, C18O 1-0 C17O 2-1, C18O 2-1 emission (from left to right) in models with thick (blue) and thin (orange) VIRaM layers. The gray dashed line in each panel indicates the location of the CO snowline. For the optically thin lines, the derivative profiles for the thick VIRaM model show a minimum near the CO snowline resulting from the significant change in the CO column density with the radius. The minimum in the derivative is less prominent in C18O 1-0 than in C17O 1-0 due to increasing optical depth, and it becomes undetectable in C18O 2-1, which is optically thick.